Dr Paul Coddington NPAC, 3-201 CST 111 College Place Syracuse University Syracuse, NY 13244

U.S.A.

Phone: (315) 443-4883 FAX: (315) 443-1973

email: paulc@npac.syr.edu

5/5/93

Dear Dr Grassberger,

I would appreciate a preprint of your paper WUB 93-03. If you

have this in electronic form I would be very grateful if you could email it to me, or if I could get it via anonymous ftp.

Yours sincerely,